OpenStudy (anonymous):

simplify. 6+2^2/52-2*6 SHOW ALL WORK

OpenStudy (anonymous):

first exponents: 6+2^2/52-2*6 6+4/52-2*6 then multiplication / division: 6+4/52-12 finally addition: 6+4/52-12 =312/52 + 4/52 - 624/52 =-308/52 =-77/13 =5.923